Paco and I messed around with the attenuation of the scope and bandwidth of the IF. We also replaced the BNC T's in the circuits with RF splitters. We saw some decent improvements to the data. The data is attached and a diagram of the experiment. [We analytically calculated the impedances to avoid any mismatch taking place]. Working on fitting the data.
We also moved around the wires so that the AG4395 is closer to the PZT.